# OpenDriveVLA: Towards End-to-end Autonomous Driving with Large Vision Language Action Model

### v1

@@ Abstract
We present OpenDriveVLA, a Vision-Language Action (VLA) model designed for end-to-end autonomous driving. OpenDriveVLA builds upon open-source pre-trained large Vision-Language Models (VLMs) to generate reliable driving actions, conditioned on 3D environmental perception, ego vehicle states, and driver commands. To bridge the modality gap between driving visual representations and language embeddings, we propose a hierarchical vision-language alignment process, projecting both 2D and 3D structured visual tokens into a unified semantic space. Besides, OpenDriveVLA models the dynamic relationships between the ego vehicle, surrounding agents, and static road elements through an autoregressive agent-env-ego interaction process, ensuring both spatially and behaviorally informed trajectory planning. Extensive experiments on the nuScenes dataset demonstrate that OpenDriveVLA achieves state-of-the-art results across open-loop trajectory planning and driving-related question-answering tasks. Qualitative analyses further illustrate OpenDriveVLA’s superior capability to follow high-level driving commands and robustly generate trajectories under challenging scenarios, highlighting its potential for next-generation end-to-end autonomous driving. We will release our code to facilitate further research in this domain.
1 Introduction
@@
End-to-end learning frameworks have emerged as a promising paradigm in autonomous driving, enabling perception, prediction, and planning to be jointly optimized within a unified neural network [54, 5]. By leveraging large-scale driving data, these models learn driving policies directly from raw sensor inputs, achieving impressive performance across diverse driving scenarios. Despite notable progress, existing end-to-end driving approaches still face critical challenges, particularly in limited generalization to long-tail scenarios, insufficient understanding of high-level semantics within complex driving scenes, and inflexible reasoning capabilities for task-driven planning [4]. In parallel, large language models (LLMs) and vision-language models (VLMs) demonstrate strong in-context learning, common sense understanding, and zero-short generalization abilities. These emergent capabilities highlight their promise for autonomous driving, especially given the critical need for robust scene understanding of diverse real-world driving conditions [28, 53]. However, directly leveraging existing VLMs for autonomous driving poses fundamental challenges. Firstly, current VLMs are predominantly optimized for static, 2D image-language tasks, leading to poor spatial reasoning performance in dynamic 3D driving environments [34, 50]. Besides, these models frequently produce hallucinated outputs—incorrect or overly confident descriptions—that significantly compromise safety in autonomous driving. Motivated by these limitations, our work answers a central question: How can we harness the emergent capabilities of VLMs to produce reliable driving actions, while mitigating hallucination risks and balancing inference speed and planning effectiveness?
To this end, we propose OpenDriveVLA, a novel vision-language action model designed explicitly for end-to-end autonomous driving. OpenDriveVLA leverages open-source pre-trained language foundation models to generate interpretable and reliable driving trajectories conditioned on multimodal inputs, including 3D environmental perception, ego vehicle state, and driver commands. OpenDriveVLA bridges vision-language understanding and trajectory generation within a unified model. It is rigorously evaluated on open-loop planning and driving-related question-answering benchmarks, achieving leading results and demonstrating superior trajectory generation and driving-scene understanding. Specifically, our contributions are summarized as follows:
@@
  • We present OpenDriveVLA, an end-to-end vision-language action model, which generates reliable driving trajectories conditioned on multimodal inputs.
  • We introduce a hierarchical vision-language feature alignment module, projecting structured 2D and 3D visual tokens into a unified semantic embedding space to facilitate language-guided trajectory generation.
  • We design an agent-env-ego interaction process to capture interactions among the ego vehicle, dynamic agents, and static map elements, significantly enhancing motion forecasting accuracy and trajectory reliability in complex traffic scenarios.
  • Extensive experiments on nuScenes dataset show that OpenDriveVLA establishes new state-of-the-art results in both open-loop planning and driving-related question answering, consistently outperforming prior LLM-based and end-to-end autonomous driving approaches.
2 Related Work
2.1 End-to-End Autonomous Driving
@@
Autonomous driving evolves through two distinct stages. Traditional approaches rely on a modular design, decomposing the system into perception [3, 40, 24], prediction [51, 30, 48], and planning [12, 37] components. While this structure ensures interpretability and allows for independent optimization, they suffer from cascading errors between stages and are not globally optimized for the final planning objective. In contrast, end-to-end autonomous driving frameworks [14, 17, 5] address this by jointly optimizing perception, prediction, and planning within a unified neural network. These models learn driving policies directly from raw sensor inputs, which improves the model’s adaptability to diverse driving conditions. However, existing end-to-end methods still face semantic reasoning bottlenecks—they struggle to fully comprehend high-level scene semantics, infer complex agent interactions, and adapt to dynamic task requirements. Moreover, their decision-making processes remain opaque, making it difficult to diagnose failure cases, especially in long-tail or unseen scenarios.
@@
(b) VLM as high-level decision-maker.
(c) OpenDriveVLA (Ours): Vision-Language-Action model as end-to-end model via integrated visual perception and language understanding.
@@
Figure 1: Three paradigms of leveraging vision language models for end-to-end autonomous driving.
Figure 2: Illustration of main training stages on OpenDriveVLA, including the hierarchical cross-modality alignment, driving instruction tuning, agent-env-ego interaction, and trajectory planning tuning.
2.2 Multi-modal Large Language Models
Large Language Models demonstrated strong emergent capabilities in in-context learning, instruction following, and reasoning [39, 33]. By training on vast amounts of Internet-scale data, these models acquire extensive world knowledge and exhibit strong adaptability across diverse tasks. Their success has also driven the rise of large VLMs, which extend these capabilities into cross-modal reasoning by integrating vision encoders with language models. State-of-the-art VLMs such as GPT-4V [31], LLaVA [26], DeepSeekVL [41], and Qwen-VL [1] demonstrate strong visual understanding and multimodal reasoning in open-domain tasks. However, these models are primarily trained on static 2D images or videos and exhibit limited spatial reasoning in dynamic 3D driving environments. Moreover, VLMs are prone to hallucinations and generally over-confident but incorrect descriptions, which pose serious risks in safety-critical planning scenarios. Recently, Vision-Language Action models have emerged to directly predict actions from visual inputs, demonstrating strong performance in robotic manipulation tasks [52, 22]. However, the application of such language-conditioned action generation in end-to-end autonomous driving remains underexplored.
2.3 Language Models in Autonomous Driving
@@
Both LLMs and VLMs have been applied to a wide range of autonomous driving tasks, including perception, scene description, synthetic data generation, and high-level decision-making [54]. Within the scope of end-to-end autonomous driving, existing works generally follow three types of design strategies. As shown in Figure 1, one line of research integrates language heads, such as captioning or question-answering modules, into driving models to enhance interpretability [8]. The second category employs language models to generate high-level driving instructions, such as directional commands or abstract maneuvers, which are subsequently interpreted by separate planning modules into low-level controls [18, 38]. This approach allows language models to influence decision-making, but retains a modular separation between reasoning and motion planning, making joint optimization challenging. The third category directly applies VLMs to predict driving actions from monocular video input [19, 44]. These methods process 2D images and generate speed and steering commands based on visual observations, without explicit modeling of the 3D spatial layout or multi-agent interactions in the driving scene. This limits their ability to reason about depth, occlusion, and agent dynamics in complex traffic environments. In this work, we aim to investigate how LLMs can unify cross-modality semantic reasoning and 3D instance-aware trajectory planning within an end-to-end autonomous driving framework.
3 OpenDriveVLA
@@
The overall architecture of OpenDriveVLA is shown in Figure LABEL:fig:title_figure, with its multi-stage training process further detailed in Figure 2. OpenDriveVLA begins with a pre-trained vision encoder, which extracts tokenized environmental representations from multi-view images (Sec. 3.1). These visual tokens are then aligned into the textual domain through cross-modal learning (Sec. 3.2). After alignment, OpenDriveVLA undergoes driving instruction tuning (Sec. 3.3), followed by agent-ego-environment interaction modeling (Sec. 3.4). Finally, OpenDriveVLA is trained end-to-end to predict the ego vehicle’s future trajectory, guided by the aligned visual-language tokens and driving instructions (Sec. 3.5).
3.1 3D Visual Environmental Perception
@@
Recent VLM-based autonomous driving methods typically rely on 2D visual encoders [34, 50], where visual token selection and attention are indirectly guided through language supervision, often in the form of question-answering tasks. While effective in open-domain vision-language applications, this design lacks explicit 3D spatial grounding and structured object-level attention, which can lead to severe hallucinations in safety-critical driving scenarios [42]. To mitigate this, OpenDriveVLA adopts a visual-centric query module, where the model first learns to focus on driving-relevant objects and map tokens through 3D vision tasks, ensuring reliable visual token proposal.
@@
3.2 Hierarchical Vision-Language Alignment
To bridge the modality gap between the extracted visual tokens and the word embedding space of a pre-trained LLM, we adopt a hierarchical vision-language feature alignment strategy. Given the visual tokens extracted from the 3D visual perception module, we introduce three token-specific projectors $\{\Phi_{\mathrm{scene}},\Phi_{\mathrm{agent}},\Phi_{\mathrm{map}}\}$. During training, each active agent query from the 3D detection and tracking task denoted as $v_{agent}^{i}$, is also matched to its corresponding ground-truth caption $\mathbf{X}_{agent}^{i}$. These captions provide detailed descriptions, including 2D appearance descriptions and 3D spatial positions. For scene and map tokens, which encode holistic spatial context and static structural properties, a sample-wise alignment is applied, where each token is matched to a scene-level caption $\mathbf{X}_{scene}$ or $\mathbf{X}_{map}$. The scene token $v_{scene}$ captures the global 2D environmental context, while the map token $v_{map}$ encodes structural elements such as lane topology, road boundaries, and drivable areas. Each of these tokens is aligned to its corresponding caption, denoted as $\mathbf{X}_{scene}$ and $\mathbf{X}_{map}$. During this stage, both the visual encoder and the pre-trained LLM remain frozen, with only the token-specific projectors being trainable. The forward alignment step is formulated as follows:
@@ 3.2 Hierarchical Vision-Language Alignment
3.3 Driving Instruction Tuning
To embed high-level driving knowledge into OpenDriveVLA and enhance its reasoning ability, we introduce a dedicated driving instruction tuning stage during training. Instead of performing explicit chain-of-thought (CoT) reasoning at inference time, which would significantly increase latency, we distill essential driving knowledge into the model through supervised instruction tuning to achieve a trade-off between inference speed and reasoning efficacy.
@@ 3.3 Driving Instruction Tuning
3.4 Agent Environment Ego Interaction
Reliable trajectory planning in autonomous driving necessitates a spatially grounded 3D representation of the environment. Beyond perception, it must also understand dynamic interactions between the ego vehicle and surrounding agents. Effective interaction modeling is essential to ensure that planned trajectories are both feasible and collision-free under real-world driving constraints. However, existing pre-trained LLMs lack an inherent inductive bias for spatial reasoning in 3D driving scenes, as they are predominantly trained on 2D vision-language and text-based datasets. To address the limitation, we introduce a conditional agent motion forecasting task, which serves as a proxy task for 3D agent-environment-ego interaction modeling, allowing the model to learn spatially grounded motion patterns. During this stage, OpenDriveVLA captures the underlying structure of multi-agent dynamics, enhancing its capability for scene-aware trajectory generation and improving decision-making in complex traffic scenarios.
@@ 3.4 Agent Environment Ego Interaction
$\operatorname*{max}\prod_{t=1}^{T}p\left(w^{i}_{t}\mid w^{i}_{1:t-1},\textbf{V}_{env},\textbf{S}_{ego}\right)$ (4)
@@
3.5 End-to-end Trajectory Planning
In this stage, OpenDriveVLA plans future driving actions as a sequence of waypoints over the next few seconds, denoted as $\mathcal{W}_{ego}=\{w_{1},w_{2},\dots,w_{T}\}$. Each waypoint $w_{t}$ represents the 2D coordinates $(x_{t},y_{t})$ of the ego vehicle at time step $t$. To enable autoregressive generation using the LLM, the waypoints are first tokenized into a sequence of discrete textual tokens: $\mathcal{T}{traj}=\mathrm{Tokenizer}(\mathcal{W}_{ego})$. The generation process is then cast as a causal sequence prediction task, where each token is predicted conditioned on the visual perception tokens $\textbf{V}_{env}$, the ego state $\textbf{S}_{ego}$, and the driving command $\textbf{X}_{dri}$:
$\hat{\mathcal{T}}_{traj}=\operatorname*{argmax}_{\textbf{T}_{traj}}\prod_{t=1}^{T}p\left(w_{t}\mid w_{1:t-1},\textbf{V}_{env},\textbf{S}_{ego},\textbf{X}_{dri}\right)$ (5)
The whole model, including the visual encoder, cross-modal projectors, and the LLM, is jointly optimized end-to-end during training. At inference, the model autoregressively generates the tokenized trajectory $\hat{\mathcal{T}}_{traj}$, which is then decoded back into numerical waypoints:
$\hat{\mathcal{W}}_{ego}=\mathrm{Decoder}(\hat{\mathcal{T}}_{traj})$ (6)
4 Experiments
4.1 Training Data
We conduct experiments on nuScenes dataset [2], following standard data split into training and validation sets. OpenDriveVLA is trained using the training set paired with corresponding QA captions, while the validation set is exclusively used for performance evaluation to ensure fair comparisons with prior works. We show the training data information for each stage in Table 1.
Vision-Language Alignment. For agent feature alignment, we post-process instance captions from [20], which provides the 2D visual description of individual objects. To further enhance spatial awareness, each object caption is augmented with its corresponding BEV coordinates, enabling the model to associate object attributes with precise spatial locations. For scene tokens, we process multi-view scene descriptions from [46], merging them into unified summaries that describe the driving environment across all camera views. For map tokens, structured language descriptions are derived from ground-truth annotations, translating map elements such as lane dividers, crosswalks, and road boundaries into descriptive text.
Driving Instruction Tuning. We adopt multiple instruction-oriented datasets derived from nuScenes to incorporate driving-specific knowledge into OpenDriveVLA. We unify several datasets into a standardized instruction-based QA format, including driving-related question-answer pairs collected from nuCaption [46], nuScenesQA [32], and nuX [8] dataset. Each QA pair is conditioned on structured environmental visual tokens and the ego vehicle state, ensuring consistency across different data sources. This multimodal instruction tuning process allows OpenDriveVLA to effectively ground language understanding into both environmental perception and scene understanding, bridging perception, reasoning, and action within the language space.
Stage 1 Stage 2 Stage 2.5 Stage 3
Data Annos 2D/3D Cap Instruct QA Trajectory Trajectory
#Samples 536k 566k 459k 28k
Model Trainable Proj Proj, LLM Proj, LLM Full Model
0.5B Size (MB) 3.1 496.9 496.9 552.6
3B Size (MB) 14.2 3099.6 3099.6 3155.3
7B Size (MB) 41.3 7654.2 7654.2 7709.9
Train 0.5B LR ($\psi_{\text{vision}}$) - - - $1\times 10^{-5}$
0.5B LR ($\{\theta_{\text{proj/LLM}}\}$) $1\times 10^{-4}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$
Epoch 1 1 1 1
Table 1: Overview of the multi-stage training pipeline in OpenDriveVLA, including data, model, and training information.
@@ 4.1 Training Data
ST-P3 [13] 1.33 2.11 2.90 2.11 0.23 0.62 1.27 0.71 - - - - - - - - - Visual
VAD [17] 0.17 0.34 0.60 0.37 0.07 0.10 0.24 0.14 - - - - - - - - - Visual
Ego-MLP [49] 0.46 0.76 1.12 0.78 0.21 0.35 0.58 0.38 - - - - - - - - - Ego
UniAD [14] 0.44 0.67 0.96 0.69 0.04 0.08 0.23 0.12 0.48 0.96 1.65 1.03 0.05 0.17 0.71 0.31 - Visual
InsightDrive [35] 0.23 0.41 0.68 0.44 0.09 0.10 0.27 0.15 0.30 0.72 1.41 0.81 0.08 0.15 0.84 0.36 - Visual
FF [11] - - - - - - - - 0.55 1.20 2.54 1.43 0.06 0.17 1.07 0.43 - LiDAR
EO [21] - - - - - - - - 0.67 1.36 2.78 1.60 0.04 0.09 0.88 0.33 - LiDAR
@@
DriveVLM [38] 0.18 0.34 0.68 0.40 0.10 0.22 0.45 0.27 - - - - - - - - Qwen-VL-7B Visual
GPT-Driver [29] 0.20 0.40 0.70 0.44 0.04 0.12 0.36 0.17 0.27 0.74 1.52 0.84 0.07 0.15 1.10 0.44 GPT-3.5 Textual
RDA-Driver [15] 0.17 0.37 0.69 0.40 0.01 0.05 0.26 0.10 0.23 0.73 1.54 0.80 0.00 0.13 0.83 0.32 LLaVa-7B Visual
EMMA [16] 0.14 0.29 0.54 0.32 - - - - - - - - - - - - Gemini [36] Visual
OpenEMMA [43] 1.45 3.21 3.76 2.81 - - - - - - - - - - - - Qwen-VL-7B Visual
DME-Driver [10] - - - - - - - - 0.45 0.91 1.58 0.98 0.05 0.28 0.55 0.29 LLaVa-7B Visual

### v2

@@ Abstract
We present OpenDriveVLA, a Vision-Language Action (VLA) model designed for end-to-end autonomous driving, built upon open-source large language models. OpenDriveVLA generates spatially-grounded driving actions by leveraging multimodal inputs, including both 2D and 3D instance-aware visual representations, ego vehicle states, and language commands. To bridge the modality gap between driving visual representations and language embeddings, we introduce a hierarchical vision-language alignment process, projecting both 2D and 3D structured visual tokens into a unified semantic space. Furthermore, we incorporate structured agent–environment–ego interaction modeling into the autoregressive decoding process, enabling the model to capture fine-grained spatial dependencies and behavior-aware dynamics critical for reliable trajectory planning. Extensive experiments on the nuScenes dataset demonstrate that OpenDriveVLA achieves state-of-the-art results across open-loop trajectory planning and driving-related question-answering tasks. Qualitative analyses further illustrate its superior capability to follow high-level driving commands and generate trajectories under challenging scenarios, highlighting its potential for next-generation end-to-end autonomous driving.
I Introduction
@@
End-to-end learning frameworks have emerged as a promising paradigm in autonomous driving, enabling perception, prediction, and planning to be jointly optimized within a unified neural network [vlmadsurvey]. They learn policies directly from sensor inputs and generalize well across varied scenarios. Despite notable progress, existing approaches still face critical challenges, including limited long-tail generalization, poor complex semantics understanding, and rigid task reasoning [chen2023e2esurvey]. Meanwhile, large language models (LLMs) and vision-language models (VLMs) exhibit strong in-context reasoning, commonsense understanding, and zero-shot generalization abilities. These capabilities are promising for driving, where robust scene understanding is crucial [datasetsurvey, zhou2024gpt4vtrafficassistantindepth]. However, directly leveraging existing VLMs for autonomous driving poses fundamental challenges. Firstly, current VLMs are predominantly optimized for static, 2D image-language tasks, leading to poor spatial reasoning performance in dynamic 3D driving environments [siglip]. Besides, instance-agnostic VLMs [liu2024surveyhallucinationlargevisionlanguage1] are prone to hallucinations, often yielding incorrect yet overconfident outputs, posing safety risks in autonomous driving. Motivated by these limitations, our work answers a central question: How can we harness the emergent capabilities of large VLMs to produce safe spatially-grounded driving actions in dynamic 3D environments, while balancing inference speed and planning effectiveness?
To enhance spatial-awareness and safety in LLM-based vision-language action model, we introduce two key designs. First, we structure the driving environment using instance-aware, hierarchical 2D and 3D visual representations to reduce the risk of instance hallucinations. Second, we incorporate agent–environment–ego interaction modeling, which is originally explicitly modeled in traditional end-to-end driving systems, as an auxiliary objective into the autoregressive LLM training pipeline. It enables the model to internalize physical feasibility and dynamic multi-agent interactions, improving robustness in safety-critical scenarios.
Built upon open-source large language models, OpenDriveVLA tightly integrates spatially-grounded multimodal reasoning and driving trajectory generation within a unified autoregressive framework. Unlike prior VLM-based methods, OpenDriveVLA leverages structured 2D and 3D instance-aware representations, ego vehicle states, and high-level commands to directly produce reliable driving actions. Extensive experiments on nuScenes benchmark demonstrate that OpenDriveVLA achieves state-of-the-art performance in both open-loop planning and vision-language reasoning tasks. Our key contributions are:
@@
  • We present OpenDriveVLA, a 3D vision-language action model for end-to-end autonomous driving that generates reliable driving trajectories by integrating hierarchical visual input, ego state, and high-level language commands.
  • We develop a multi-stage training strategy that aligns structured 2D and 3D visual features into a unified semantic space, enabling naive VLMs to generate spatially-grounded actions in complex driving scenarios.
  • We introduce implicit agent–environment–ego interaction modeling into autoregressive LLM-based VLA training as an auxiliary task, enabling the model to learn behaviorally grounded and safety-aware driving actions.
II Related Work
II-A End-to-End Autonomous Driving
@@
Autonomous driving (AD) evolves through two distinct stages. Traditional approaches rely on a modular design, decomposing the system into perception [bevformer], prediction [pred1], and planning [planning1] components. While this structure ensures interpretability and allows for independent optimization, they suffer from cascading errors between stages and are not globally optimized for the final planning objective. In contrast, end-to-end autonomous driving frameworks [hu2023_uniad] address this by jointly optimizing perception, prediction, and planning within a unified neural network. These models learn driving policies directly from raw sensor inputs, which improves the model’s adaptability to diverse driving conditions. More recent approaches introduce diffusion models [diffusiondrive] and unified scene representations [jia2025drivetransformer] to further enhance the effectiveness and robostness. However, existing end-to-end methods still face semantic reasoning bottlenecks, as they struggle to fully comprehend high-level scene semantics, infer complex agent interactions, and adapt to dynamic task requirements. Moreover, their decision-making processes remain opaque, making it difficult to diagnose failure cases, especially in long-tail or unseen scenarios.
II-B Large Vision Language Models
Large Language Models demonstrated strong emergent capabilities in in-context learning, instruction following, and reasoning [touvron2023llamaopenefficientfoundation, qwen2.5]. By training on vast amounts of Internet-scale data, these models acquire extensive world knowledge and exhibit strong adaptability across diverse tasks. Their success has also driven the rise of large VLMs, which extend these capabilities into cross-modal reasoning by integrating vision encoders with language models. State-of-the-art VLMs such as GPT-4V [openai2024gpt4technicalreport], LLaVA [liu2024llavanext], and Qwen-VL [Qwen-VL] demonstrate strong visual understanding and multimodal reasoning in open-domain tasks. However, these models are primarily trained on static 2D images or videos and exhibit limited spatial reasoning in dynamic 3D driving environments. Moreover, VLMs are prone to hallucinations and generally over-confident but incorrect descriptions, which pose serious risks in safety-critical planning scenarios. Recently, Vision-Language Action models have emerged to directly predict actions from visual inputs, demonstrating strong performance in robotic manipulation tasks [kim24openvla]. Currently, the application of such language-conditioned end-to-end action generation in autonomous driving remains underexplored. Yet, these methods are mostly limited to static setups and lack driving-specific 3D spatial design.
@@ II-B Large Vision Language Models
(b) VLM as high-level driving decision-maker.
(c) Native 2D VLM for end-to-end driving.
(d) 3D spatial-aware driving VLA (ours).
@@
Figure 1: Taxonomy of vision-language model applications in end-to-end autonomous driving.
II-C Vision Language Models in Autonomous Driving
@@
VLMs have been applied to various autonomous driving tasks, including perception, scene description, synthetic data generation, and high-level decision-making [vlmadsurvey]. These efforts aim to enhance interpretability, data efficiency, and instruction-following capabilities in driving models. We categorize recent works into 4 paradigms, as illustrated in Figure 1. One line of research in Fig.1 (a) integrates language heads, such as captioning or question-answering modules, into driving models to enhance the interpretability [hintad-nux]. The second category in Fig.1 (b) employs vision language models to generate high-level driving instructions, such as directional commands or abstract maneuvers, which are subsequently interpreted by separate planning modules into low-level controls [jiang2024senna, tian2024drivevlm, wang2023drivemlm]. It’s also usually formed as a fast-slow dual system. This design allows VLMs to make independent semantic reasoning, but retains a separate module for end-to-end driving planning, making joint optimization challenging. The third line in Fig.1 (c) applies native VLMs with 2D visual tokens to produce driving actions, and optionally scene captions or QA responses [adapt, drivegpt4]. These methods [leapAD, 10655906, fu2025orionholisticendtoendautonomous] process 2D images without explicit modeling of the instance, 3D spatial layout, and inter-agent interactions in the driving scene. It limits their spatial reasoning ability and understanding of agent dynamics in complex traffic environments. Recent studies [favero2024multimodalhallucinationcontrolvisual2] further indicate that such instance-agnostic approaches are more prone to hallucinate, often producing overconfident or semantically inconsistent text. In this work, we investigate how to extend 2D VLMs by explicitly modeling 3D instance-aware and spatial-aware scene representations into an end-to-end autonomous driving framework, as shown in Fig.1(d). Notably, we focus on fully differentiable end-to-end models in this work, while LLM-based agentic driving systems, such as [wang2024omnidrive, sima2023drivelm], fall outside the scope of our study.
III OpenDriveVLA
@@
The overall architecture of OpenDriveVLA is shown in Figure LABEL:fig:title_figure, with its multi-stage training process further detailed in Figure 2. OpenDriveVLA uses a pre-trained vision encoder to extract tokenized environmental representations from multi-view images. These visual tokens are then aligned into the textual domain through cross-modal learning. After alignment, it undergoes driving instruction tuning, followed by agent-ego-environment interaction modeling. Finally, OpenDriveVLA is trained end-to-end to predict the ego vehicle’s future trajectory, guided by the aligned visual-language tokens and driving instructions.
Figure 2: Illustration of main training stages on OpenDriveVLA. Stage 1: Hierarchical Feature Alignment. Stage 2: Driving Instruction Tuning. Stage 2.5: Agent-Env-Ego Interaction Modeling. Stage 3: Trajectory Planning Tuning.
III-A 3D Visual Environmental Perception
@@
Recent VLM-based autonomous driving methods typically rely on pretrained 2D visual encoders [siglip], where visual token selection and attention are indirectly guided through language supervision. While effective in open-domain vision-language applications, this design lacks explicit 3D spatial grounding and structured instance-level attention, which can lead to severe hallucinations in safety-critical driving scenarios [drivebench]. To mitigate this, OpenDriveVLA adopts a visual-centric query module, where the model first learns to focus on driving-relevant objects and map tokens through 3D vision tasks, ensuring reliable visual token proposal.
@@
III-B Stage 1 - Hierarchical Vision-Language Alignment
To bridge the modality gap between the extracted visual tokens and the word embedding space of a pre-trained LLM, we adopt a hierarchical vision-language feature alignment strategy. Given the visual tokens extracted from the 3D visual perception module, we introduce three token-specific projectors $\{\Phi_{\mathrm{scene}},\Phi_{\mathrm{agent}},\Phi_{\mathrm{map}}\}$. During training, each active agent query from the 3D detection and tracking task denoted as $v_{agent}^{i}$, is also matched to its corresponding ground-truth caption $\mathbf{X}_{agent}^{i}$. These captions provide detailed descriptions, including 2D appearance descriptions and 3D spatial positions. For scene and map tokens, which encode holistic spatial context and static structural properties, a sample-wise alignment is applied, where each token is matched to a scene-level caption $\mathbf{X}_{scene}$ or $\mathbf{X}_{map}$. The scene token $v_{scene}$ captures the global 2D environmental context, while the map token $v_{map}$ encodes structural elements such as lane topology, road boundaries, and drivable areas. Each of these tokens is aligned to its corresponding caption, denoted as $\mathbf{X}_{scene}$ and $\mathbf{X}_{map}$. During this stage, both the visual encoder and LLM remain frozen to preserve pretrained semantics, with only the token-specific projectors being trainable. The forward alignment step is formulated as follows:
@@ III-B Stage 1 - Hierarchical Vision-Language Alignment
III-C Stage 2 - Driving Instruction Tuning
We distill high-level driving knowledge into the model via supervised instruction tuning, enabling it to internalize semantic reasoning patterns during training. This avoids costly chain-of-thought (CoT) reasoning at inference time and balances planning efficacy with runtime efficiency.
@@ III-C Stage 2 - Driving Instruction Tuning
III-D Stage 2.5 - Agent Environment Ego Interaction
Reliable trajectory planning in autonomous driving necessitates a spatially grounded 3D representation of the environment. Beyond perception, it must also understand dynamic interactions between the ego vehicle and surrounding agents. Effective interaction modeling is essential to ensure that planned trajectories are both feasible and collision-free under real-world driving constraints. However, existing pre-trained LLMs lack an inherent inductive bias for spatial reasoning in 3D driving scenes, as they are predominantly trained on 2D vision-language and text-based datasets. We introduce a conditional agent trajectory forecasting task as an auxiliary objective, encouraging the model to learn spatially grounded interaction priors. During this stage, OpenDriveVLA captures the underlying structure of multi-agent dynamics, enhancing its capability for scene-aware trajectory generation and improving decision-making in complex traffic scenarios.
@@
$\mathrm{max}\prod_{t=1}^{T}p\left(w^{i}_{t}\mid w^{i}_{1:t-1},\textbf{V}_{env},\textbf{S}_{ego},\Phi_{\mathrm{agent}}(v_{agent}^{i})\right)$ (4)
@@
ST-P3 [stp3] 1.33 2.11 2.90 2.11 0.23 0.62 1.27 0.71 - - - - - - - - - Visual
VAD [vad] 0.17 0.34 0.60 0.37 0.07 0.10 0.24 0.14 - - - - - - - - - Visual
Ego-MLP [ego-mlp] 0.46 0.76 1.12 0.78 0.21 0.35 0.58 0.38 - - - - - - - - - Ego
UniAD [hu2023_uniad] 0.44 0.67 0.96 0.69 0.04 0.08 0.23 0.12 0.48 0.96 1.65 1.03 0.05 0.17 0.71 0.31 - Visual
InsightDrive [song2025insightdriveinsightscenerepresentation] 0.23 0.41 0.68 0.44 0.09 0.10 0.27 0.15 0.30 0.72 1.41 0.81 0.08 0.15 0.84 0.36 - Visual
FF [planning1] - - - - - - - - 0.55 1.20 2.54 1.43 0.06 0.17 1.07 0.43 - LiDAR
EO [EO] - - - - - - - - 0.67 1.36 2.78 1.60 0.04 0.09 0.88 0.33 - LiDAR
@@
GPVL [GPVL] 0.21 0.39 0.69 0.43 0.07 0.09 0.27 0.14 - - - - - - - - BERT Textual
DriveVLM [tian2024drivevlm] 0.18 0.34 0.68 0.40 0.10 0.22 0.45 0.27 - - - - - - - - Qwen-VL-7B Visual
GPT-Driver [mao2024gptdriver] 0.20 0.40 0.70 0.44 0.04 0.12 0.36 0.17 0.27 0.74 1.52 0.84 0.07 0.15 1.10 0.44 GPT-3.5 Textual
RDA-Driver [RDA-Driver] 0.17 0.37 0.69 0.40 0.01 0.05 0.26 0.10 0.23 0.73 1.54 0.80 0.00 0.13 0.83 0.32 LLaVa-7B Visual
OminiDrive [wang2024omnidrive] 0.14 0.29 0.55 0.33 0.00 0.13 0.78 0.30 - - - - - - - - LLaVA-7B Visual
EMMA [hwang2024emmaendtoendmultimodalmodel] 0.14 0.29 0.54 0.32 - - - - - - - - - - - - Gemini Visual
OpenEMMA [xing2025openemmaopensourcemultimodalmodel] 1.45 3.21 3.76 2.81 - - - - - - - - - - - - Qwen-VL-7B Visual
DME-Driver [dmedriver] - - - - - - - - 0.45 0.91 1.58 0.98 0.05 0.28 0.55 0.29 LLaVa-7B Visual